OpenStudy (anonymous):

NEED HELP!! WILL MEDAL!! GEOMETRY What is the ratio of the area of square ABCF to the area of rectangle ABDE?

OpenStudy (janu16):

Do you have picture?

OpenStudy (dayakar):

draw a rough figure

OpenStudy (deer_slayer13):

d

OpenStudy (janu16):

What?

OpenStudy (deer_slayer13):

D(14,14)EFC(11,10)AB(5,2)y

OpenStudy (janu16):

Plot carefully, or use properties of slope You 'll find F(5,8) So the ratio is [ (5 x 5) / ( 5+15) x 5 ] = 5/20 =1/4

OpenStudy (anonymous):

So its 1/4?

OpenStudy (janu16):

Yea